OpenStudy (anonymous):

Which of the following has a solution of 3? 6 times the reciprocal of 1/2 5 ÷ 5/3 4/5 of 7 1/2 2/5 • 2/15

OpenStudy (dangerousjesse):

\(\large\frac{2}{1}\times 6=?\)

OpenStudy (08surya):

1

OpenStudy (dangerousjesse):

No, not one :)

OpenStudy (anonymous):

6 times the reciprocal of 1/2

OpenStudy (08surya):

yes sorry

OpenStudy (anonymous):

12

OpenStudy (anonymous):

it should be reciprocal of 2 not 1/2

OpenStudy (anonymous):

No it says 1/2

OpenStudy (dangerousjesse):

The reciprocal of 1/2 is definitely \(\large\frac{1}{2}=\frac{2}{1}=2\)

OpenStudy (anonymous):

What's the answer

OpenStudy (dangerousjesse):

You were right in saying 12 :)

OpenStudy (anonymous):

Oh coo

OpenStudy (dangerousjesse):

\(\Huge\frac{5}{\frac{5}{3}}\)

OpenStudy (dangerousjesse):

Can you solve that? :)

OpenStudy (anonymous):

Uhh u don't really know how to solve that

OpenStudy (anonymous):

I sorry

OpenStudy (dangerousjesse):

No problem haha. What's \(\large5\div 3\)?

OpenStudy (anonymous):

15

OpenStudy (dangerousjesse):

Not \(\large 5\times 3\), \(\large 5\div 3\). :)

OpenStudy (anonymous):

1.67

OpenStudy (anonymous):

I don't know

OpenStudy (dangerousjesse):

Yes, and \(\large 5\div 1.67\)?

OpenStudy (anonymous):

2990142

OpenStudy (dangerousjesse):

2.99, yes, which rounds to..?

OpenStudy (anonymous):

3.

OpenStudy (dangerousjesse):

Yep :)

OpenStudy (anonymous):

But I don't know the choice I should put @ dangerousJesse